# Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star forming galaxies at z5.2 with the N2CLS

### v1

@@ 3.2 Properties of N2GN_1_13
The total IR luminosity (integrated between 8 and 1000 m), SFR, stellar mass , dust mass and dust spectral index have been derived in B25 using multi-wavelength SED fitting.
The dust mass has been obtained using Draine (2003) frequency-dependent dust absorption coefficient, , renormalized as indicated by Draine et al. (2014), i.e. m2 kg-1 as reference. The panchromatic optical to radio SEDs are modeled with the MAGPHYS and SED3FIT codes, in their original and high- versions (da Cunha et al. 2008, 2015; Battisti et al. 2020; Berta et al. 2013), as well as with CIGALE (Burgarella et al. 2005; Noll et al. 2009; Boquien et al. 2019). The spectral index is obtained using a modified black-body fit to the far-IR to mm data. The effect of the cosmic microwave background is taken into account as described by da Cunha et al. (2013).
The stellar mass has been obtained using the same assumptions for CIGALE and the high- versions of MAGPHYS and SED3FIT, i.e. Bruzual & Charlot (2003, BC03) stellar populations with a Chabrier (2003) IMF and the Charlot & Fall (2000) two-components attenuation law. B25 presents the comparison of stellar masses obtained using the different codes, and shows a very good agreement, but for few outliers lying at high z. Unfortunately, the sources considered in this paper are among the outliers. For example, considering the 7 optically-dark sources at z5.2 (see Table 3), the ratio between the high- versions of MAGPHYS or SED3FIT and CIGALE stellar mass determination ranges from 1.4 to 9.2, with an average ratio of 5.2.
@@
Excluding any AGN contamination, the high flux observed in the long-wavelength NIRCam/JWST F444W filter (0.7-1.3 Jy) supports by itself the very large stellar masses, i.e. M1011-1012 M⊙. Such very high stellar masses goes with very large extinction.
For N2GN_1_13, the stellar masses from SED3FIT and CIGALE are log(M⋆)=11.68 and log(M⋆)=11.03, respectively. In the following, and to be conservative, we adopt the lower stellar mass value obtained using CIGALE. To be consistent, all the other parameters (e.g. SFR, dust mass) are taken from the CIGALE SED fitting.
@@
Following Berta et al. (2023) (see also Dunne et al. 2022), we adopt a conversion factor = 4.0 M⊙ (K km s-1 pc2)-1, including the contribution of helium. The luminosity of the CO lines detected by NOEMA are transformed into
adopting the average ratios given by Carilli & Walter (2013) for SMGs, , and , which give (and also , see Sect. 4) consistent with our measurement. We obtain = (2.30.6)1010 M⊙ and (3.10.9)1010 M⊙ from the CO(5-4) and CO(4-3) line luminosity, respectively. We adopt the molecular gas mass from the CO(4-3) which is detected with a higher SNR and with a measurable FWHM (Table 1).
We can now derive the depletion time of the molecular gas in each galaxy using the molecular gas mass and the SFR:
This is the time it would take for a galaxy to consume the entirety of its molecular gas reservoir, if it formed stars at the rate derived from far-IR and sub-mm data in a single event (not considering any mass return to the ISM from stellar winds and supernovae).
We also finally compute the gas-to-dust mass ratio GDR=MMol/Mdust.
@@
N2GN_1_13 has a very high SFR and is thus in the process of very efficient stellar mass build-up. It has a specific star formation rate log(sSFR)=-8.18. Compared to the main sequence of star formation, N2GN_1_13 lies above, with SFR = log(SFR/SFRMS)= 0.55 (if we use the parametrization of the MS from Popesso et al. 2023) and SFR = 0.29 (if we use the parametrization of Speagle et al. 2014). The former exhibits the well-known flattening at large stellar masses, whereas the latter does not.
@@
N2GN_1_13 has a very red continuum with ABmag. As shown in B25, the optical photometry reveals a rest-frame UV excess, which is reproduced by adding an unextinguished young (ages between 10 and 100 Myr) simple stellar population to the SED modeling. As can be seen on Fig. 7, it has an extended morphology, in contrast with the so-called little red dots, LRD (e.g., Labbé et al. 2023; Greene et al. 2024; Matthee et al. 2024; Pérez-González et al. 2024). The source is very red simply because it is dusty with i) a very high Av=3.960.13 mag, indicating a large dust content and ii) rest-frame dust emission detected by NOEMA at both 324 and 574 m. Furthermore, this source cannot be classified as a broad-line AGN, as the line does not exhibit broadening. It has no X-ray counterpart and no rest-frame mid-IR data are available to verify the possible presence of an AGN-torus component contributing to its mid-IR SED.
@@
The very high SFR relative to the available gas reservoir results in a short depletion time of 43 Myr, indicative of a brief (100 Myr) starburst phase555Adopting = 1 M⊙ (K km s-1 pc2)-1 (e.g., Riechers et al. 2020) would give a depletion time 4 times lower, making this source even more extreme. The 3 upper limit on the CI(1-0) line gives M1010M⊙ and 120 Myr, assuming a conversion factor = 17 M⊙ (K km s-1 pc2)-1 from Dunne et al. (2022).. Such a low points to an extreme case of star formation. Assuming no gas accretion or replenishment, and modeling the gas consumption as an exponential decay , the galaxy’s SFR would drop to 10% of its initial value within approximately 100 Myr, effectively marking the quenching of star formation.
@@
With a dust mass of 7.70.8108 M⊙, and given its stellar mass, N2GN_1_13 lies near the maximal theoretical dust yield expected from the combined contributions of supernovae (assuming no reverse-shock destruction) and AGB stars (see Fig. 5 and Witstok et al. 2023). However, the short gas depletion time suggests that dust production from AGB stars, whose contribution operates on longer timescales, is likely minimal. This implies that additional dust formation mechanisms, such as rapid grain growth in the ISM, may be required to account for the observed dust mass.
@@
Our observed gas-to-dust ratio leads to a galaxy metallicity consistent with solar, (e.g., Li et al. 2019). However, this estimate should be interpreted with caution, as our adopted values for and are themselves based on the assumption of solar-like metallicity.
@@
With a CO(5–4)/CO(4–3) flux ratio of 1.37, N2GN_1_13 closely matches the ratio observed in the galaxy CGG-z4.b at (Brinch et al. 2025), and falls within the upper range of values reported in the literature (see the Fig. 5 in Brinch et al. 2025). This elevated ratio may suggest the presence of high gas temperatures, densities, and/or pressures.
@@
N2GN_1_13 is thus a massive, extended, and highly obscured galaxy caught in a short-lived yet extreme starburst phase at z5.2. Its high stellar mass, red optical continuum, and gas reservoir point to rapid evolution and imminent quenching.
@@
4 Gas masses, gas-to-dust ratios, gas depletion times for the 3 brightest DSFGs in the overdense environment
In addition to N2GN_1_13, two other galaxies in the overdense environment have CO measurements: GN10 and HDF850.1 (see the compilation in Riechers et al. 2020, R20). We use the same recipes as in Sect. 3 to compute gas masses, GDR and gas depletion times for these two galaxies and use the other source properties (e.g., SFR) as derived from the SED fitting with CIGALE. The results are given in Table 2.
We use the CO(2-1) line measurements for HDF850.1 and GN10 and the conversion factor from Carilli & Walter (2013), which gives MMol consistent with R20 if we would consider =1 rather than =4, and assuming a mean gravitational lensing magnification factor =1.6 for HDF850.1 as in R20. If we further correct for (which is equal to 0.057 m2 kg-1 in R20 versus 0.047 here), we would obtain GDR=46 and 31 w.r.t. 65 and 130 in R20, and =39 and 34 Myr w.r.t to 70 and 40 Myr in R20, for GN10 and HDF850.1, respectively.
The large difference in GDR for HDF850.1 arises from a factor 3.7 difference in Mdust (assuming the same ), while the discrepancy in for GN10 is mainly due to a factor 1.6 in SFR. Notice that we have the same spectral index as in previous studies (R20, Walter et al. 2012).
We derive gas mass fraction fgas = MMol / (MMol+M⋆) of 0.6 for HDF850.1 and GN10, which is typical of gas fraction of z5 sub-millimeter galaxies (Dudzevičiūtė
et al. 2020). For N2GN_1_13, the gas mass fraction is 3 times lower but remains consistent with certain z5 SMGs reported in the literature (e.g. Zavala et al. 2022). However, we warn that such measurements have to be taken with caution.
For instance, Sun et al. (2024) reported a gas fraction of 0.18 for HDF850.1, which is significantly lower than the value obtained in this study. This discrepancy primarily arises from the considerably lower gas mass reported in Neri et al. (2014), attributed to the adoption of a much lower (this is also the case in Zavala et al. (2022) where =1.0 M⊙ (K km s-1 pc2)-1).
Tacconi et al. (2020) give updated scaling relations that links M⋆, redshift and distance from the main sequence of star forming galaxies (from Speagle et al. 2014) to molecular gas depletion timescale () and thus molecular gas mass (M). We use these relations to compute and M (see Table 4) and obtain = 206, 142, 161 Myr and M=1.41011, 4.71010, 2.41011 M⊙ for N2GN_1_13, HDF850.1 (delensed) and GN10, respectively. While this compares quite well to the results given in Table 2 for HDF850.1 and GN10 (10% level), M (and thus ) is overestimated by a factor 4.3 for N2GN_1_13. If we were using higher J transitions to get the molecular mass for HDF850.1 and GN10, we would have obtained the same discrepancy (i.e. molecular gas mass of 3.1 and 5.41010M⊙ using Jup=5 rather than 12 and 251010 M⊙). This shows how these determinations have to be taken with caution.
@@ 5 Bright dusty galaxies in the overdense environment
N2GN_1_01 | GN10, S2 | 12:36:33.44 | 62:14:08.7 | 5.303s | Optically-dark | |
N2GN_1_06 | HDF850.1 | 12:36:51.99 | 62:12:25.7 | 5.185s | Optically-dark | |
N2GN_1_13 | GN32 | 12:36:58.61 | 62:14:51.1 | 5.182s | Optically-dark | |
N2GN_1_23 | GN15, S3 | 12:36:56.54 | 62:12:07.5 | 5.179s | Optically-dark | |
N2GN_1_61 | P-40 | 12:36:57.48 | 62:16:54.4 | 5.201s | Optically-dark | |
N2GN_1_20 | AzGN4 | 12:35:50.22 | 62:10:42.4 | 5.330.32p | 5.16 | Optically-dark |
N2GN_1_43 | 12:37:11.95 | 62:11:19.7 | 5.870.38p | 5.31 | Optically-dark | |
@@
†: The HDF850, GN, AzGN, HRG14, S and P- nomenclatures were defined by Hughes et al. (1998), Pope et al. (2005), Perera et al. (2008), (Finkelstein et al. 2015), Xiao et al. (2024), Greve et al. (2008).
‡ s: spectroscopic redshift; p: photometric redshift from B25.
h photometric redshift from New-Hyperz
@@ 5.1 N2CLS galaxies at similar redshifts
In addition to N2GN_1_01 (GN10), N2GN_1_06 (HDF850.1) and N2GN_1_13 discussed above, we have 7 more galaxies with redshifts compatible with the overdense structure (see Table 3).
Two are optically-dark N2CLS galaxies with spectroscopic redshift from JWST: N2GN_1_23 (also known as S3 in (Xiao et al. 2024) and N2GN_1_61.
For the five additional galaxies with previously reported photometric redshifts of (see B25 and references therein), independent photometric redshifts were also derived using the enhanced JWST photometric coverage (when available) and the SED-fitting code New-Hyperz. This is a modified version of the public code Hyperz originally developed by Bolzonella et al. (2000)666https://people.osupytheas.fr/~rpello/newhyperz/. The procedure and its validation is discussed in Appendix B. Results are shown in Table 3. Our new photometric redshifts are within one-sigma uncertainty of the literature based , for all but N2GN_1_26 (2 ). The new photometric redshifts support their association with the overdensity.
Among these five N2CLS galaxies, two are optically dark (N2GN_1_20 and N1_1_43), while the remaining three are optically bright (N2GN_1_26, N2GN_1_33, N2GN_1_57).
Details on each source are provided in Appendix D.
On the 10 sources (5 with and 5 with ), N2GN_1_01, 06 and 23 are already identified to belong to the overdensity in Sun et al. (2024); Herard-Demanche et al. (2025); Xiao et al. (2024) ; N2GN_1_13 and N2GN_1_61 fall into their area of study but were not previously identified. ; N2GN_1_33 and 43 are in the FRESCO and JADES footprint, respectively, but were not reported ; N2GN_1_57 lies at the border of the JADES footprint and N2GN_1_20, 26 are further away.

### v2

@@ 3.2 Properties of N2GN_1_13
The total IR luminosity (; integrated between 8 and 1000 m), SFR, stellar mass ), dust mass (), and dust spectral index () have been derived in B25 using multi-wavelength SED fitting.
The dust mass has been obtained using draine2003 frequency-dependent dust absorption coefficient (), re-normalised as indicated by draine2014, i.e. m2 kg-1 as reference. The panchromatic optical to radio SEDs are modelled with the MAGPHYS and SED3FIT codes, in their original and high- versions (dacunha2008; dacunha2015; battisti2020; berta2013), as well as with CIGALE (burgarella2005; noll2009; boquien2019). The spectral index is obtained using a modified black-body fit to the far-IR to millimetre data. The effect of the cosmic microwave background is taken into account as described by dacunha2013.
The stellar mass has been obtained using the same assumptions for CIGALE and the high- versions of MAGPHYS and SED3FIT, i.e. bc03 stellar populations with a chabrier2003 IMF and the charlot2000 two-components attenuation law. B25 presents the comparison of stellar masses obtained using the different codes, and shows a very good agreement, but for few outliers lying at high z. Unfortunately, the sources considered in this paper are among the outliers. For example, considering the 6 optically dark sources at z5.2 (see Table 3), the ratio between the high- versions of MAGPHYS or SED3FIT and CIGALE stellar mass determination ranges from 1.4 to 9.2, with an average ratio of 5.8.
@@
Excluding any active galactic nucleus (AGN) contamination, the high flux observed in the long-wavelength NIRCam/JWST F444W filter (0.7-1.3 Jy) supports by itself the very large stellar masses, i.e. M1011-1012 M⊙. Such very high stellar masses go with high extinction.
For N2GN_1_13, the stellar masses from SED3FIT and CIGALE are log(M⋆)=11.68 and log(M⋆)=11.03, respectively. In the following, and to be conservative, we adopt the lower stellar mass values obtained using CIGALE. To be consistent, all the other parameters (e.g. SFR, dust mass) are taken from the CIGALE SED fitting.
@@
Following berta2023, we adopted a conversion factor = 4.0 M⊙ (K km s-1 pc2)-1, including the contribution of helium. We transformed the luminosity of the CO lines detected by NOEMA into
by adopting the average ratios given by Carilli2013 for sub-millimetre galaxies (SMGs), , and , which give (and also ; see Sect. 4) consistent with our measurement. We obtain = (2.30.6)1010 M⊙ and (3.10.9)1010 M⊙ from the CO(5-4) and CO(4-3) line luminosity, respectively. We adopted the molecular gas mass from the CO(4-3), which is detected with a higher S/N and with a measurable full width at half maximum (Table 1).
We could then derive the depletion time of the molecular gas using the molecular gas mass and the SFR:
This is the time it would take for a galaxy to consume the entirety of its molecular gas reservoir, if it formed stars at the rate derived from far-IR and sub-millimetre data in a single event (not considering any mass return to the interstellar medium (ISM) from stellar winds and supernovae).
We also finally computed the gas-to-dust mass ratio (GDR) as MMol/Mdust.
@@
N2GN_1_13 is a very massive galaxy (1.071011 M⊙). Using the same survey volume as FRESCO (Xiao2024), such a high mass requires a high fraction of the baryons (30%) to be converted into stars (; as discussed in Sect. 6). This is significantly higher than what is commonly observed in lower-z galaxies (Xiao2024; gentile2024).
@@
N2GN_1_13 has a very red continuum with ABmag. As shown in B25, the optical photometry reveals a rest-frame UV excess, which is reproduced by adding an unextinguished young (ages between 10 and 100 Myr) simple stellar population to the SED modelling. As can be seen in Fig. 7, it has an extended morphology. The source is very red simply because it is dusty with (i) a very high Av=3.960.13 mag, indicating a large dust content and (ii) rest-frame dust emission detected by NOEMA at both 324 and 574 m. Furthermore, this source cannot be classified as a broad-line AGN, as the H line does not show any additional broadening beyond what is expected from morphological and kinematic effects. It has no X-ray counterpart and no rest-frame mid-IR data are available to verify the possible presence of an AGN-torus component contributing to its mid-IR SED.
@@
The very high SFR relative to the available gas reservoir results in a short depletion time of 43 Myr, indicative of a brief (100 Myr) starburst phase777Adopting = 1 M⊙ (K km s-1 pc2)-1 (e.g. Riechers2020) would give a depletion time 4 times lower, making this source even more extreme.. Such a low points to an extreme case of star formation. Assuming no gas accretion or replenishment, and modelling the gas consumption as an exponential decay , the galaxy’s SFR would drop to 10% of its initial value within approximately 100 Myr, effectively marking the quenching of star formation.
@@
With a dust mass of (7.70.8)108 M⊙, and given its stellar mass, N2GN_1_13 lies near the maximal theoretical dust yield expected from the combined contributions of supernovae (assuming no reverse-shock destruction) and asymptotic giant branch (AGB) stars (see Fig. 6 and witstok2023). However, the short gas depletion time suggests that dust production from AGB stars, whose contribution operates on longer timescales, is likely minimal. This implies that additional dust formation mechanisms, such as rapid grain growth in the ISM, may be required to account for the observed dust mass.
@@
Our observed gas-to-dust ratio leads to a galaxy metallicity consistent with solar, (e.g. li2019). However, this estimate should be interpreted with caution, as our adopted values for and are themselves based on the assumption of solar-like metallicity.
@@
With a CO(5–4)/CO(4–3) flux ratio of 1.37, N2GN_1_13 closely matches the ratio observed in the galaxy CGG-z4.b at (Brinch2025), and falls within the upper range of values reported in the literature (see Fig. 5 in Brinch2025). This elevated ratio suggests the presence of high gas temperatures, densities, and/or pressures.
N2GN_1_13 has a very high SFR and is thus in the process of very efficient stellar mass build-up. Compared to the star-forming main sequence (SFMS), N2GN_1_13 lies significantly above the relation if we adopt the parametrisation of popesso2023, with SFR = log(SFR/SFRMS)= 0.550.02 (for a scatter of 0.09 dex). Using the parametrisation of speagle2014, however, the offset is smaller, SFR = 0.290.04 (for a scatter of 0.2 dex). These SFMS relations were calibrated with pre-JWST data, and it is now well established that massive galaxies at z5 can remain undetected in HST observations. As a result, the SFMS at the high-mass end (M M⊙) is not necessarily well constrained. Considering the specific star-formation rate (sSFR), N2GN_1_13 has log(sSFR)=-8.180.13, which is close to the empirical limit of log(sSFR)-8.05 for main-sequence (MS) galaxies reported by caputi2017, based on star-forming galaxies selected from optical and near-IR surveys.
Taken together, these results suggest that N2GN_1_13 is broadly consistent with, though slightly above, the SFMS, within the current uncertainties at the high-mass end.
@@
N2GN_1_13 is thus a massive, extended, and highly obscured galaxy caught in a short-lived starburst phase at z5.2. Its high stellar mass, red optical continuum, and gas reservoir point to rapid evolution and imminent quenching.
@@
4 Gas masses, gas-to-dust ratios, and gas depletion times for the three brightest DSFGs in the overdense environment
In addition to N2GN_1_13, two other galaxies in the overdense environment have CO measurements: GN10 and HDF850.1 (see the compilation in Riechers2020, R20). We used the recipes described in Sect. 3 to compute gas masses, GDR and gas depletion times for these two galaxies and use the other source properties (e.g. SFR) as derived from the SED fitting with CIGALE. The results are given in Table 2.
We used the CO(2-1) line measurements for HDF850.1 and GN10 and the conversion factor from Carilli2013, which gives MMol consistent with R20 if we consider =1 rather than =4, and assuming a mean gravitational lensing magnification factor =1.6 for HDF850.1 as in R20. If we further correct for (which is equal to 0.057 m2 kg-1 in R20 versus 0.047 here), we would obtain GDR=46 and 31 w.r.t. 65 and 130 in R20, and =39 and 34 Myr w.r.t to 70 and 40 Myr in R20, for GN10 and HDF850.1, respectively.
The large difference in the GDR for HDF850.1 arises from a factor of 3.7 difference in Mdust (assuming the same ), while the discrepancy in for GN10 is mainly due to a factor of 1.6 difference in the SFR. Notice that we have the same spectral index as in previous studies (R20, walter2012).
We derive a gas mass fraction (fgas = MMol / (MMol+M⋆)) of 0.6 for HDF850.1 and GN10, which is typical of gas fraction of z5 SMGs (S2CLS_UDS2020). For N2GN_1_13, the gas mass fraction is 3 times lower but remains consistent with certain z5 SMGs reported in the literature (e.g. Zavala2022). However, we warn that such measurements have to be taken with caution.
For instance, Sun2024 reported a gas fraction of 0.18 for HDF850.1, which is significantly lower than the value obtained in this study. This discrepancy primarily arises from the considerably lower gas mass reported in Neri2014, attributed to the adoption of a much lower (this is also the case in Zavala2022 where =1.0 M⊙ (K km s-1 pc2)-1).
tacconi2020 give updated scaling relations that links M⋆, redshift and distance from the MS of star-forming galaxies (from speagle2014) to molecular gas depletion timescale () and thus molecular gas mass (M). We used these relations to compute and M (see Table 4) and obtain = 206, 142, 161 Myr and M=1.41011, 5.11010, 2.61011 M⊙ for N2GN_1_13, HDF850.1 (de-lensed) and GN10, respectively. While this compares quite well to the results given in Table 2 for HDF850.1 and GN10 (5% level), M (and thus ) is overestimated by a factor of 4.5 for N2GN_1_13. If we used higher J transitions to get the molecular mass for HDF850.1 and GN10, we would have obtained the same discrepancy (i.e. molecular gas mass of 3.1 and 5.41010M⊙ using Jup=5 rather than 12 and 251010 M⊙). This shows how these determinations have to be taken with caution.
@@ 5 Bright dusty galaxies in the overdense environment
†: The HDF850, GN, HRG14, S and P- nomenclatures were defined by hughes1998, pope2005, finkelstein2015, Xiao2024, greve2008.
‡ s: spectroscopic redshift and p: photometric redshift, from B25 ; h photometric redshift from New-Hyperz.
@@
| N2GN_1_01 | GN10, S2 | 12:36:33.44 | 62:14:08.7 | 5.303s | Optically dark | |
| N2GN_1_06 | HDF850.1 | 12:36:51.99 | 62:12:25.7 | 5.185s | Optically dark | |
| N2GN_1_13 | GN32 | 12:36:58.61 | 62:14:51.1 | 5.182s | Optically dark | |
| N2GN_1_23 | GN15, S3 | 12:36:56.54 | 62:12:07.5 | 5.179s | Optically dark | |
| N2GN_1_61 | P-40 | 12:36:57.48 | 62:16:54.4 | 5.201s | Optically dark | |
| N2GN_1_43 | 12:37:11.95 | 62:11:19.7 | 5.870.38p | 5.31 | Optically dark | |
@@ 5.1 N2CLS galaxies at similar redshifts
In addition to N2GN_1_01 (GN10), N2GN_1_06 (HDF850.1), and N2GN_1_13, we have five more galaxies with redshifts and position compatible with the overdense structure101010N2CLS found an other dusty galaxy at z5.3, which is farther away (see Appendix D). (see Table 3). Two are optically dark N2CLS galaxies with spectroscopic redshift from JWST: N2GN_1_23 (also known as S3 in (Xiao2024) and N2GN_1_61. For the three additional galaxies with previously reported photometric redshifts of (see B25 and references therein), independent photometric redshifts were also derived using the enhanced JWST photometric coverage (when available) and the SED-fitting code New-Hyperz. This is a modified version of the public code Hyperz originally developed by Bolzonella_2000111111https://people.osupytheas.fr/~rpello/newhyperz/. The procedure and its validation is discussed in Appendix B. Results are shown in Table 3. Our new photometric redshifts are within one-sigma uncertainty of the literature based , for all but N2GN_1_26 (2 ). The new photometric redshifts therefore support their possible association with the z5.1-5.4 overdensity. While photometric redshifts provide the best constraints currently available, spectroscopic confirmation will be required to secure their membership. For example, a photometric-redshift-selected galaxy, N2GN_1_33, initially estimated at z5.3 (B25) was later confirmed with JWST spectroscopy to lie at z=1.78, illustrating the uncertainties inherent in photometric redshift estimates. In the following, we therefore clearly distinguish results for galaxies with spectroscopic and photometric redshifts. Among the three N2CLS galaxies with zphot, one is optically dark (N1_1_43), while the remaining two are optically bright (N2GN_1_26 and N2GN_1_57). Details on each of the five sources are provided in Appendix C.
Of the eight sources (five with and three with ), N2GN_1_01, 06, and 23 belong to the overdensity according to Sun2024, Herard-Demanche2024, and Xiao2024; N2GN_1_13 and N2GN_1_61 fall into their area of study but were not identified. N2GN_1_43 is in the JADES footprint but was not reported. N2GN_1_57 lies at the border of the JADES footprint and N2GN_1_26 is farther away.
† corrected from =2.5.